For most flying creatures, the length of the wing varies inversly as the number of
wingbeats per second. A dragonfly has a winglengh of 1.6 inches and a wingspeed of 45
times per second. If a bee has a wingspeed of 200 beats per second, what is the length
of the bee's wing? (Simply work the problem through; you do not need a 5-step solution.)
My work so far:
45 Divided by 1.6 = 28.125
But what do I do now? Please help me!!! I'm stuck!! Thanks! :)

Found 3 solutions by Alan3354, jrfrunner, edjones:
Answer by Alan3354(69443)

For most flying creatures, the length of the wing varies inversly as the number of
wingbeats per second. A dragonfly has a winglengh of 1.6 inches and a wingspeed of 45
times per second. If a bee has a wingspeed of 200 beats per second, what is the length
of the bee's wing? (Simply work the problem through; you do not need a 5-step solution.)
My work so far:
45 Divided by 1.6 = 28.125
--------------------
w = length of wing
f = frequency
If they vary inversely, f = k/w k is some constant
45 = k/1.6
k = 72
---------
For the bee:
200 = 72/w
w = 72/200 = 0.36 inch

Answer by jrfrunner(365)
Answer by edjones(8007)

Let the 2 inversely related variables be x and y.
xy=k [a constant] True for all inversely related things.
1.6*45=72
200x=72
x=0.36" length of a bee's wing.
.
Ed